You just decided to begin saving for retirement. You will make deposits of $1,000 per month into a retirement account that earns 8.00% p.a. The first deposit is made today and the last deposit will be made when you retire exactly 30 years from today. You will begin to make withdrawals from the account the first month after you retire. If you plan to live an addition 25 years and leave $800,000 to your heirs, you will be able to withdraw $_____. (Note: you make 300 total monthly withdrawals from your retirement account.)

please use finance calculator to solve
Answer #1

Using financial calculator

Input:

PMT = -1000

I/Y = 8/12

N = 30*12 = 360

BGN Mode

Solve for FV as 1,500,295.18

Using financial calculator

Input:

PV = -1,500,295.18

N = 25*12 = 300

I/Y = 8/12

FV = 800000

Solve for PMT as 10,738.33

Amount that can be withdrawn = $10,738.33